# Underage Detection through a Multi-Task and MultiAge Approach for Screening Minors in Unconstrained Imagery

### v1

@@ 5 Experiments and Discussion
This section presents an empirical evaluation of our proposal. We begin by reproducing a baseline age estimator (Section 5.1), then explore training on the unified dataset (Section 5.2). Section 5.3 investigates under-18 detection under various training settings, while Section 5.4 extends these results to other age thresholds and explores architectural variations.
@@ 5.1 Age Estimation Baseline
To establish a performance baseline, we start reproducing the state of the art of age estimation by running the “FaRL + MLP” model of [15] in several deployment conditions. Unlike the original approach, we omit the pre-training of MLP on the IMDB-WIKI dataset to simplify the training process and isolate the impact of our contributions.
In order to illustrate the effect of the mismatch between training and deployment conditions, we set up an experiment where the model is trained with different datasets and applied to different scenarios. Table 2 reports the Mean Absolute Error (MAE) metric, which refers to the absolute value of the deviation between estimated age and true age, averaged over the respective test dataset.
@@
Table 2: Mean Absolute Error (MAE) (in years) obtained in the baseline setting of the Facial-Age-Estimation-Benchmark of [15] with a “FaRL + MPL” model. Each row is for a training set (benchmark), and each column is for a test dataset (holdout set in the case of intra-dataset evaluation). All rows are averaged over a five-fold cross validation, except CACD2000 and CLAP2016 where only one split is used.
Table 3 shows the differences with respect to the values reported by Paplhám and Franc [15]. As we have mentioned before, note that their experiments, involved pre-traininig of the MLP. This is probably the reason why we measure worse accuracy for the rows “CACD2000” and “MORPH”, i.e. models trained with these two datasets, one with label noise and one with limited variability, and evaluated on a different dataset. Even without pre-training of the MLP, in all, we reach a similar accuracy as Paplhám and Franc [15], since in some cases we measured lower errors (negative differences in Table 3).
@@
Table 3: Difference of the Mean Absolute Error (MAE) given in Table 2 with respect to the values given in [15, Table 6, line “FaRL + MPL”]. For some entries our accuracy is worse (positive difference); for others it is better (negative numbers).
@@ 5.2 Age estimation with an overall training set
Apart from the individual benchmarks, we also employ the overall benchmark as described in Section 3.1.1, which is constructed as a joint training set together with consistent holdout sets for validation and testing.
@@
Table 4: Mean Absolute Error (MAE) (in years) obtained with the overall training sets. The rows show the overall benchmarks “Overall” and “Overall without CACD” trained with the union of six, and five datasets, respectively, and evaluated on the test set, averaged over five folds. The mark (0) is for future reference.
We consider two variants, the overall training set with and without the CACD data. The test errors obtained with the different test datasets are listed in Table 4 and represented graphically in Figure 6. It is shown that the model trained with the Overall training set achieves low age-regression errors on all datasets. Indeed, the accuracy of the model trained with the Overall training set achieves accuracies close to the respective intra-dataset accuracies, indicating that the capacity of the “FaRL + MLP” model is sufficient to adapt to the data distributions of the individual benchmarks simultaneously, providing a “one fits all” model.
@@ 5.3 Under-18 detection
To assess how different training settings and training strategies affect under-18 detection, we evaluate multiple variants of our pipeline.
  1. (i)
    Data cleaning: Whether the training data was cleaned or raw.
  2. (ii)
    Dataset inclusion: Whether the noisy CACD2000 dataset was included or not.
  3. (iii)
    Sampling strategy: Whether age-balanced resampling was applied in training.
  4. (iv)
    Model architecture: Whether a dedicated binary under-18 head was added to the model.
For the binary discrimination task of under-18 detection, under-18 and over-18 are the positive and the negative class, respectively, and the confusion matrix is given by $((\mathrm{TP},\mathrm{FN}),(\mathrm{FP},\mathrm{TN}))$, where F/T stand for false/true and P/N for positive negative. The most important error measure is the false over-age detection rate, i.e. the fraction of underage subjects (positives) that are incorrectly classified as adults (negatives), also known as the false negative rate $\mathrm{FNR}=\mathrm{FN}/(\mathrm{TP}+\mathrm{FN})$. This value should be low with high priority. In detection-error tradeoff (DET) curves, the false adult detection rate $\mathrm{FNR}$ is plotted against the false positive rate $\mathrm{FPR}=\mathrm{FP}/(\mathrm{TN}+\mathrm{FP})$ or false under-age detection rate, parametrized by the threshold applied to the confidence score (or the age estimate used to separate under-age from adults).
A complementary view on the confusion matrix is given in terms of precision and recall. Recall $=\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})=1-\mathrm{FNR}$ measures whether all children were detected correctly. It should be high (close to one) with high priority. Precision $=\mathrm{TP}/(\mathrm{TP}+\mathrm{FP})$] measures whether all detected children are really children. It should also be high (close to 1), with lower priority than recall. The $F\beta$ scores interpolate between Recall and Precision. The formula $F\beta=(1+\beta^{2})/(1/\mathrm{precision}+\beta^{2}/\mathrm{recall})$ reduces to the harmonic mean in the case of F1 and gives higher priority to recall in the case of F2.
@@ 5.3.1 Impact of the training set and training scheme
We consider the Overall training set of 3.1.1, with and without the cleaning of label noise (Section 3.1.2), as well as a variant where we remove the CACD2000 dataset from the training. We add a dedicated non-adult (under 18) detector to the model and train it with binary cross-entropy loss, optionally with the age-balanced resampling (Section 4.2).
Figure 7 shows the detection error tradeoff (DET) curve for under/over 18-years discrimination measured on the ASORES-39k (restricted) test (Section 3.2.1). Black dots with dotted lines (“age only”) correspond to the row marked as (0) in Table 4, where the age estimate is used for under-18 discrimination. In the other curves, the output of the under-18 head of the network is used, and the curves are parameterized by the threshold applied.
Figure 7: Under/over 18-years discrimination accuracy obtained under different training schemes (cln: cleaned or not; CACD: with or without the CACD part of the training set; res: using resampling or not) and with or without a dedicated under/over 18 discriminator. The default working points, marked with thick dots, are “age estimate = 18” for the age estimator and over age probability = sigmoid(0) = 0.5 for the binary under-18 discriminator. The shaded area marks the desired range of operation with a low false over-age detection rate of $10^{-2}$. Three configurations of interest are marked for future reference (0, A, B, C).
Note that the default working points are strongly biased towards false-over age detections because the training data contains much more adults than underage subjects (see age distribution in Figure 2, curve “Total”). In order to achieve low false under-age detection rates, as required for age verification, the threshold has to be adjusted significantly. For example, to achieve a false-adult detection rate as low as $10^{-2}$, the threshold of the plain age estimator has to be raised from 18 to 32 years.
The other curves of Figure 7 show how the under-18 discrimination accuracy is improved by:
  1. 1.
    Adding a binary discriminator as a second head to the model, which is trained for under-18 discrimination, in parallel (curves with “+ under-18” in the legend).
  2. 2.
    Modifying the training set by cleaning (Section 3.1.2) and/or by removing the noisy CACD2000 dataset from the training.
  3. 3.
    Age-equalized resampling during training (“res = yes”, described in Sec. 4.2).
Three setups for the training perform equally well at the working point of false-adult detection rate = $10^{-2}$ (the gray region in Figure 7), the FaRL + MLP architecture with age estimation and under-18 heads, trained with the cleaned “Overall” training set (A), additionally with resampling (B), and omitting the the CACD data in the training (C). Out of these, the last one, model C, outperforms the others in a broad range of the DET curve, in particular at the point of equal error rate (diagonal of the plot).
@@ 5.3.2 Additional underage detectors
Figure 8: Under/over 18-years discrimination accuracy. Variation points: additional age-verification discriminators (under-12, 15, 18, 21); omitting CACD2000 from the training, resampling and reweighting. All configurations use the cleaned training sets (cln=yes in Figure 7).
To improve the under-18 discrimination, we explore the addition of binary detectors by exploring the next two variation points:
  1. 1.
    Further over-age discriminators (i.e., non-adult detectors at age levels 12, 15, 18, 21 years) to increase the focus on the underage region.
  2. 2.
    Configure the $\alpha$ parameter. The alpha parameter applied in the loss function to correct for the remaining class imbalance (Section 4.2).
Figure 8 shows DET curves comparing models with only an under-18 head to those with additional discriminators on ASORES-39k. The first four curves in Figure 8, are taken from the previous Figure 7. The curves show the under-18 discrimination accuracy of models that incorporate further under-age discriminators at age levels 12, 15, and 18 years. Training the model with resampling and the $\alpha$ parameter (light orange and light green) turns out worse than using just the resampling (dark orange and dark green). Best accuracy for the under-18 discrimination task is achieved by model D, i.e. using the training set consisting only of AFAD, AgeDB, CLAP2016, Morph Album 2, and UTKFace (but not CACD), trained with age-balanced resampling (Sec. 4.2) and with further under-age discriminators (but without further tuning of the $\alpha$ parameter).
@@
ASORES-39k (restricted test) ASWIFT-20k (wild test)
@@
yes 0.583 0.734 0.869 0.990 0.464 0.630 0.804 0.985
Table 5: Quantitative comparison of under-18 detection accuracy. The discriminator threshold is fixed such that the false-adult detection rate is 0.01 on the restricted overall test (shaded area in Figure 8), i.e. Recall = 0.99. Precision (Pr) measures if detected non-adults are really adults. Recall (Re) measures whether all non-adults have been detected as such. F1 and F2 scores interpolate between both. For each measure (column), the two experiments with the best values are highlighted.
Pose-45 expression image-conditions
Heads CACD res $\alpha$ Pr F1 F2 Re Pr F1 F2 Re Pr F1 F2 Re
age
with no no 0.934 0.958 0.973 0.984 0.423 0.591 0.776 0.982 0.233 0.376 0.596 0.976 (0)
age + u-18 with no no 0.955 0.972 0.983 0.990 0.528 0.689 0.843 0.992 0.321 0.483 0.691 0.969 (A)
yes no 0.955 0.975 0.987 0.995 0.514 0.677 0.835 0.990 0.301 0.459 0.671 0.967 (B)
without yes no 0.954 0.974 0.987 0.995 0.531 0.690 0.841 0.984 0.309 0.467 0.675 0.959 (C)
age + 12, 15, 18, 21 with no no 0.952 0.972 0.984 0.993 0.514 0.677 0.836 0.991 0.312 0.473 0.684 0.973
yes no 0.945 0.968 0.982 0.992 0.482 0.649 0.818 0.991 0.279 0.434 0.650 0.972
yes 0.942 0.965 0.980 0.990 0.473 0.640 0.812 0.988 0.272 0.426 0.643 0.974
without yes no 0.955 0.975 0.988 0.996 0.527 0.688 0.842 0.990 0.311 0.472 0.685 0.981 (D)
yes 0.959 0.975 0.984 0.991 0.514 0.677 0.836 0.992 0.301 0.459 0.671 0.969
Table 6: Break down of the Wild test into its three main parts, otherwise like Table 5.
In Table 5, we collect quantitative measures at the discriminator threshold defined above for the restricted overall test (same as in Figure 8). The columns referring to the ASORES-39k restricted test reflect the ranking of the curves in the gray region in Figure 8, and the best accuracy is achieved in the second-last row model D. It achieves an F2 score of 0.875 on the Restricted Overall Test and 0.813 on the ASWIFT-20k wild test, both at a fixed recall of 0.99. This represents an absolute F2 improvement of +0.074 and +0.071, respectively, over the baseline age estimator. The columns for the Wild Test in the same table, as well as Table 6, prove that the good accuracy of this model carries nicely over to unconstrained conditions. In fact, model D is the only model that is able to maintain the Recall value of 0.99 in the Wild test.
@@ 5.4 Performance across age thresholds and effects on age estimation
We now evaluate whether adding multi-task heads benefits detection at the other relevant age thresholds (12, 15, and 21). In addition, we examine how these multi-task heads influence age regression performance. For the quantification of the regression error we switch from the MAE performance metric, which was used in Section 5.1 for the sake of comparability, to the root mean square error (RMSE). The RMSE penalizes the worst estimation errors more strongly than the MAE and, thus, is the stronger criterion for the reliability of a regressor because a low RMSE requires the estimation error not only to be low on average but also to be consistent in the sense that the estimates are reasonably good for all data points.
@@ 5.4.1 Impact of the training set, training scheme, and heads
In Table 7, the quantitative analysis is extended to the underage detection tasks, with age levels, 12, 15, 18, and 21 years, and to the age estimation task, in the form of the F2 score and the RMSE, respectively.
Test ASORES-39k (restricted test) ASWIFT-20k (wild test)
F2 F2 F2 F2 RMSE F2 F2 F2 F2 RMSE
Heads CACD res $T\!=\!12$ $T\!=\!15$ $T\!=\!18$ $T\!=\!21$ (years) $T\!=\!12$ $T\!=\!15$ $T\!=\!18$ $T\!=\!21$ (years)
age only with no 0.748 0.752 0.801 0.846 5.733 0.666 0.689 0.742 0.830 6.827 (0)
yes 0.818 0.775 0.812 0.854 5.671 0.751 0.718 0.757 0.831 6.798
without no 0.905 0.831 0.833 0.853 5.687 0.905 0.757 0.764 0.838 6.577
yes 0.843 0.815 0.833 0.854 5.726 0.781 0.748 0.770 0.836 6.606
age + under-18 with no 0.846 0.800 0.869 0.861 5.499 0.778 0.728 0.817 0.842 6.526 (A)
yes 0.846 0.783 0.868 0.857 5.655 0.769 0.712 0.803 0.837 6.789 (B)
without no 0.847 0.771 0.848 0.841 5.675 0.786 0.695 0.785 0.832 6.567
yes 0.926 0.847 0.870 0.858 5.648 0.910 0.784 0.807 0.838 6.567 (C)
age + 12,15,18,21 with no 0.960 0.929 0.865 0.885 5.422 0.940 0.897 0.811 0.857 6.498
yes 0.939 0.826 0.863 0.875 5.704 0.846 0.776 0.787 0.847 6.836
without no 0.961 0.947 0.864 0.869 5.591 0.952 0.918 0.796 0.843 6.368
yes 0.959 0.938 0.875 0.875 5.602 0.933 0.909 0.813 0.851 6.497 (D)
@@
Table 7: Quantitative measurement of the underage detection (L2 score) at the other age levels $T=12,15,18,21$, together with the age regression error measures (RMSE) for the Restricted Overall Test and the Wild Test. The variation points are the inclusion of heads for under-age discrimination, whether to include CACD in the (overall) training set, and whether to use the age-equalized resampling. All experiments use the weight-sharing MLP (denoted $(512)^{2}$ in Table 8 below). For each measure (column), the top three experiments are highlighted.
@@
Several general observations can be made. Higher F2 scores can be achieved for lower age thresholds (under-12-years F2 scores up to 0.96) than for higher age ones (under-21-years F2 score below 0.89). This is a property of the data and reflects the fact that at young ages facial features change more rapidly with age than at older ages; It is easier to tell apart under-12 years old from the rest of the population than, e.g., under 21-years old from the rest. Further, as expected, the ASWIFT-20k test is harder than the ASORES-39k test, regarding all underage discrimination measures and also the age-estimation error measures.
@@
The age-regression error improves with respect to the baseline (model 0) in the first line as soon as one or several underage detectors are added. This is probably due to (a) an effect of synergy of the related tasks, but could (b) also be the regularizing effect of an additional task that avoids overfitting.
Obviously, adding dedicated under-age detectors to the model is also an advantage over using the age estimate for these discrimination tasks. Thus, the under-age discriminators are beneficial for all tasks and in both the restricted and the wild test case.
@@
Within each setting for the heads, the trend is that the combination of removing CACD and age-equalized sampling leads to superior accuracy. In summary, to this point, we have developed two techniques to improve on age estimation and non-adult detection with respect to the baseline (top line in Table 7):
  • Add under-age discriminators has a direct benefit for under-age detection, but is also beneficial for age estimation (synergy of related tasks, regularizing effect).
  • Curate the training data by removing noisy age labels and age-equalization.
@@
These results confirm that multi-threshold detection heads - model D - improves classification performance across age boundaries and also enhances age regression accuracy.

### v2

@@ 5 Experiments and Discussion
This section presents an empirical evaluation. First, we reproduce a baseline age estimator (Section 5.1), then explore training on the unified dataset (Section 5.2). Section 5.3 investigates under-18 detection under various training settings, while Section 5.4 extends the results to other thresholds and explores architectural variations.
@@ 5.1 Age Estimation Baseline
To establish a performance baseline, we reproduce the state of the art in age estimation by running the “FaRL + MLP” model of [Paplham_CVPR2024_benchmark] under several deployment conditions. Unlike the original approach, we omit the pre-training of MLP on the IMDB-WIKI dataset to simplify the training process and isolate the impact of our contributions.
To illustrate the effect of the mismatch between training and deployment conditions, we set up an experiment in which the model is trained on different datasets and applied to other scenarios. Table 2 reports the Mean Absolute Error (MAE) metric, which refers to the absolute value of the deviation between estimated age and true age, averaged over the respective test dataset.
@@
Table 2: Mean Absolute Error (MAE) (in years) obtained in the baseline setting of the Facial-Age-Estimation-Benchmark of [Paplham_CVPR2024_benchmark] with a “FaRL + MLP” model. Each row corresponds to a training set (benchmark), and each column to a test dataset (a holdout set in the case of intra-dataset evaluation). All rows are averaged over a five-fold cross-validation, except CACD2000 and CLAP2016, where only one split is used.
Table 3 shows the differences with respect to the values reported by Paplham_CVPR2024_benchmark. As mentioned above, their experiments involved pre-training the MLP, whereas ours did not. This is probably the reason why we measure worse accuracy in the rows “CACD2000” and “MORPH”, i.e. for models trained with these two datasets, one with label noise and one with limited variability. In some cases, we measured lower errors (negative differences in Table 3); thus, in general, we reach a similar accuracy as Paplham_CVPR2024_benchmark, even without pre-training the MLP.
@@
Table 3: Difference of the Mean Absolute Error (MAE) given in Table 2 with respect to the values given in [Paplham_CVPR2024_benchmark, Table 6, line “FaRL + MLP”]. For some entries, our accuracy is worse (positive difference); for others, it is better (negative numbers).
@@ 5.2 Age estimation with an overall training set
Apart from the individual benchmarks, we also employ the overall benchmark as described in Section 3.1, which is constructed as a joint training set together with consistent holdout sets for validation and testing.
@@
Table 4: Mean Absolute Error (MAE) (in years) obtained with the overall training sets. The rows show the overall benchmarks “Overall” and “Overall without CACD” trained with the union of six and five datasets, respectively, and evaluated on the test set, averaged over five folds. The mark (0) is for future reference.
We consider two variants, the overall training set with and without the CACD data. The test errors obtained with the different test datasets are listed in Table 4 and represented graphically in Figure 6. The model trained on the Overall training set achieves low age-regression errors across all datasets. Indeed, the model trained on the Overall training set achieves accuracies close to the respective intra-dataset accuracies, indicating that the capacity of the “FaRL + MLP” model is sufficient to simultaneously adapt to the data distributions of the individual benchmarks, providing a “one fits all” model.
@@ 5.3 Under-18 detection
To assess how different training settings and training strategies affect under-18 detection, we evaluate four variants of our pipeline: data cleaning (whether the training data was cleaned or raw); dataset inclusion (if the noisy CACD2000 dataset was included or not); sampling strategy (whether age-balanced resampling was applied in training); and model architecture (if a dedicated binary under-18 head was added).
For the binary discrimination task of under-18 detection, under-18 and over-18 are the positive and the negative class, respectively, and the confusion matrix is given by $((\mathrm{TP},\mathrm{FN}),(\mathrm{FP},\mathrm{TN}))$, where F/T stands for false/true and P/N for positive/negative. The most important error measure is the false over-age detection rate, i.e. the fraction of underage subjects (positives) that are incorrectly classified as adults (negatives), also known as the false negative rate $\mathrm{FNR}=\mathrm{FN}/(\mathrm{TP}+\mathrm{FN})$. This value should be low with high priority. In detection-error tradeoff (DET) curves, the false adult detection rate $\mathrm{FNR}$ is plotted against the false positive rate $\mathrm{FPR}=\mathrm{FP}/(\mathrm{TN}+\mathrm{FP})$ or false underage detection rate, parametrized by the threshold applied to the confidence score (or the age estimate used to separate underage from adults).
Recall and precision offer a complementary view on the confusion matrix. Recall $\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})=1-\mathrm{FNR}$ measures whether all minors were detected correctly. It should be high with high priority. Precision $\mathrm{TP}/(\mathrm{TP}+\mathrm{FP})$ measures whether all detected minors are really minors. It should also be high, but with lower priority. The $F\beta$ scores interpolate between Recall and Precision. The formula $F\beta=(1+\beta^{2})/(1/\mathrm{precision}+\beta^{2}/\mathrm{recall})$ reduces to the harmonic mean for F1 and gives higher priority to recall in the case of F2, as it is adequate for the use case of CSEM detection.
@@ 5.3.1 Impact of the training set and training scheme
We consider the Overall training set of 3.1, with and without label noise cleaning (Section 3.1), as well as a variant that excludes the CACD2000 dataset from the training. We add a dedicated underage (under-18) detector to the model and train it with binary cross-entropy loss, optionally with the age-balanced resampling (Section 4.2).
Figure 7 shows the detection error tradeoff (DET) curve for under/over 18-year discrimination measured on the ASORES-39k (restricted) test (Section 3.2.1). Black dots with dotted lines (“age only”) correspond to the row marked as (0) in Table 4, where the age estimate is used for under-18 discrimination. In the other curves, the output of the under-18 head of the network is used, and the curves are parameterized by the threshold applied.
Figure 7: Under/over 18-years discrimination accuracy obtained under different training schemes (cln: cleaned or not; CACD: with or without the CACD part of the training set; res: using resampling according to Eq. 7 or not and with or without a dedicated under/over 18 discriminator). Common parameters: $\alpha=1,\gamma=0,G=0$. The default working points, marked with thick dots, are “age estimate = 18” for the age estimator and over age probability = 0.5 for the binary under-18 discriminator. Shaded area: desired range of operation, false over-age detection rate $\approx 10^{-2}$. The models (0, A, B, C) are marked for future reference.
Note that the default working points are strongly biased towards false-over-age detections because the training data contains much more adults than underage subjects (see age distribution in Figure 2(b), curve “Total”). To achieve low false underage detection rates, as required in practice, the threshold must be adjusted significantly. For example, to achieve a false-adult detection rate as low as $10^{-2}$, the threshold of the plain age estimator has to be raised from 18 to 32 years.
The other curves of Figure 7 show how the under-18 discrimination accuracy is improved by (1) adding a binary discriminator as a second head to the model, which is trained for under-18 discrimination (curves with “+ under-18” in the legend); (2) modifying the training set by cleaning (end of Section 3.1) and/or by removing the noisy CACD2000 dataset from the training; (3)applying age-equalized resampling during training (“res = yes”, described in Sec. 4.2). Three setups for the training perform equally well at the working point of false-adult detection rate = $10^{-2}$ (the gray region in Figure 7), the FaRL + MLP architecture with age estimation and under-18 heads, trained with the cleaned “Overall” training set (A), additionally with resampling (B), and omitting the CACD data in the training (C). Out of these, the last one, model C, outperforms the others in a broad range of the DET curve, in particular at the point of equal error rate (diagonal of the plot).
@@ 5.3.2 Additional underage detectors
Figure 8: Under/over 18-year discrimination accuracy. Variation points: use CACD2000 in the training or not; resampling (res) according to Eq. 7 or not; prediction head setup (Heads); reweighting (alpha) according to Eq. 6 or $\alpha_{y}=1$. Common parameters: cleaned training sets (cln=yes in Figure 7), $\gamma=0,G=0$.
To improve the under-18 discrimination, we explore the addition of binary detectors by exploring the next two variation points: (1) the addition of further over-age discriminators (i.e., underage detectors at age levels 12, 15, 18, 21 years) to increase the focus on the underage region and (2) the configuration of the $\alpha$ parameter applied in the loss function to correct for the remaining class imbalance (Section 4.2).
Figure 8 shows DET curves comparing models with only an under-18 head to those with additional discriminators on ASORES-39k. The first four curves, in Figure 8, are taken from the previous Figure 7. The curves show the under-18 discrimination accuracy of models that incorporate further underage discriminators at age levels 12, 15, and 18 years. Training the model with resampling and the $\alpha$ parameter (light orange and light green) turns out worse than using just the resampling (dark orange and dark green). Best accuracy for the under-18 discrimination task is achieved by model D, i.e. using the training set consisting only of AFAD, AgeDB, CLAP2016, MORPH Album 2, and UTKFace (but not CACD), trained with age-balanced resampling (Sec. 4.2) and with further underage discriminators (but without further tuning of the $\alpha$ parameter).
@@
ASORES-39k (restricted) ASWIFT-20k (unconstrained)
@@
without yes yes 0.583 0.734 0.869 0.990 0.464 0.630 0.804 0.985
Table 5: Quantitative comparison of under-18 detection accuracy of the models shown in Figure 8. The discriminator threshold is fixed such that the false-adult detection rate is 0.01 on the restricted overall test (shaded area in Figure 8), i.e. Recall = 0.99. Precision (Pr) measures if detected minors are truly minors. Recall (Re) measures whether all minors have been correctly detected. F1 and F2 scores lie between the two measures. For each measure (column), the three experiments with the best values are highlighted.
In Table 5, we present quantitative measures at the discriminator threshold defined above for the restricted overall test (same as in Figure 8). The columns referring to the ASORES-39k restricted test reflect the ranking of the curves in the gray region in Figure 8, and the best accuracy is achieved in the second-last row model D. It achieves an F2 score of 0.875 on the Restricted Overall Test and 0.813 on the ASWIFT-20k wild test, both at a fixed recall of 0.99. This represents an absolute improvement of +0.074 and +0.071 in F2, respectively, over the baseline age estimator model 0. The columns for the ASWIFT-20k test in the same table prove that the reasonable accuracy of this model carries nicely over to unconstrained conditions. In fact, model D is the only model that can maintain the Recall value of 0.99 in the Wild test.
@@ 5.4 Performance across age thresholds and effects on age estimation
We now evaluate whether adding multi-task heads benefits detection at the other relevant age thresholds (12, 15, and 21). In addition, we examine how these multi-task heads influence age regression performance.
@@ 5.4.1 Impact of the training set, training scheme, and heads
In Table 6, the quantitative analysis is extended to the underage detection tasks, with age levels 12, 15, 18, and 21 years, quantified by the F2 score.
ASORES-39k (restricted test) ASWIFT-20k (unconstrained)
F2 F2 F2 F2 MAE F2 F2 F2 F2 MAE
Heads CACD res T=12T\!=\!12 T=15T\!=\!15 T=18T\!=\!18 T=21T\!=\!21 (years) T=12T\!=\!12 T=15T\!=\!15 T=18T\!=\!18 T=21T\!=\!21 (years)
age only
with no 0.748 0.752 0.801 0.846 4.175 0.666 0.689 0.742 0.830 4.818 (0)
yes 0.818 0.775 0.812 0.854 4.145 0.751 0.718 0.757 0.831 4.836
without no 0.905 0.831 0.833 0.853 4.034 0.905 0.757 0.764 0.838 4.441
yes 0.843 0.815 0.833 0.854 4.174 0.781 0.748 0.770 0.836 4.688
age + under-18
with no 0.846 0.800 0.869 0.861 4.022 0.778 0.728 0.817 0.842 4.603 (A)
yes 0.846 0.783 0.868 0.857 4.178 0.769 0.712 0.803 0.837 4.838 (B)
without no 0.847 0.771 0.848 0.841 4.052 0.786 0.695 0.785 0.832 4.464
yes 0.926 0.847 0.870 0.858 4.098 0.910 0.784 0.807 0.838 4.618 (C)
age + 12,15,18,21
with no 0.960 0.929 0.865 0.885 3.919 0.940 0.897 0.811 0.857 4.501
yes 0.939 0.826 0.863 0.875 4.173 0.846 0.776 0.787 0.847 4.807
without no 0.961 0.947 0.864 0.869 3.977 0.952 0.918 0.796 0.843 4.313
yes 0.959 0.938 0.875 0.875 4.062 0.933 0.909 0.813 0.851 4.468 (D)
@@
Table 6: Quantitative measurement of the underage detection (F2 score) at the other age levels $T=12,15,18,21$ and the age regression error measure MAE for the Restricted Overall Test and the Wild Test. The variation points are the inclusion of heads for underage discrimination, whether to include CACD in the (overall) training set, and whether to use the age-equalized resampling. Experiments use the weight-sharing MLP (denoted WS in Table 7 below). For each measure (column), the top three experiments are highlighted.
@@
Several general observations can be made. Higher F2 scores can be achieved for lower age thresholds (under-12-year F2 scores up to 0.96) than for higher age thresholds (under-21-year F2 scores below 0.89). This is a property of the data and reflects the fact that facial features change more rapidly with age at young ages than at older ages; It is easier to distinguish under-12-year-olds from the rest of the population than, e.g., under-21-year-olds from the rest. Further, as expected, the ASWIFT-20k test is harder than the ASORES-39k test, regarding both underage discrimination and age estimation.
@@
The age-regression error improves relative to the baseline (model 0) in the first line as soon as one or several underage detectors are added. This is related to three effects: (a) a synergistic effect between the related tasks, (b) a regularizing effect of an additional task avoiding overfitting, and (c) a dedicated underage detector is an advantage over using the age estimate for the underage detection tasks. The underage detectors are beneficial for all tasks and in both the restricted and the wild test cases.
@@
Among the settings for the heads, the trend is that the combination of removing CACD and age-equalized sampling leads to superior accuracy. In summary, to this point, we have developed two techniques to improve on age estimation and underage detection with respect to the baseline (mark (0) in Table 6): (1) Adding underage discriminators has a direct benefit for underage detection, but is also beneficial for age estimation (synergy of related tasks, regularizing effect); (2) Curate the training data by removing noisy age labels and age-equalization.
@@
These results confirm that multi-threshold detection heads - model D - improve classification performance across age boundaries and enhance age regression accuracy.